For a Saturday matinee, adult tickets cost $5.50 and kids under 12 pay only $3.00. If 110 tickets are sold for a total of $480, how many of the tickets were adult tickets and how many were sold to kids under 12?
Answer by fractalier(6550)

Call the numbers of adult and kids' tickets, a and k.
Then we write
a + k = 110 and the value equation
5.5a + 3k = 480
Now let's multiply the top one by 3 and subtract...we get
5.5a + 3k = 480
-(3a + 3k = 330)
------------------
2.5a = 150
a = 60 adult tix which means there were
k = 50 kid's tix